Question Help Suppose that a customer is purchasing a car. He conducts an experiment in which he puts 10 gallons of gas in the car and drives it until it runs out of gas. He conducts this experiment 15 times on each car and records the number of miles driven. Full data set Car 1 235 235 208 208 239 239 216 216 208 208 297 297 285 285 159 159 298 298 251 251 164 164 307 307 283 283 305 305 263 263 Car 2 215 215 205 205 232 232 221 221 246 246 248 248 241 241 272 272 243 243 298 298 280 280 246 246 259 259 294 294 261 261 Describe each data​ set, that is determine the​ shape, center, and spread. Sample mean for Car 1 x overbar equals x= nothing mi​ / 10 gal ​(Type an integer or decimal rounded to one decimal place as​ needed.) Sample mean for Car 2 x overbar equals x= nothing mi​ / 10 gal ​(Type an integer or decimal rounded to one decimal place as​ needed.) Median for Car 1 M equals = nothing mi​ / 10 gal ​(Type an integer or decimal rounded to one decimal place as​ needed.) Median for Car 2 M equals = nothing mi​ / 10 gal ​(Type an integer or decimal rounded to one decimal place as​ needed.) Range for Car 1 R equals = nothing mi​ / 10 gal ​(Type an integer or decimal rounded to one decimal place as​ needed.) Range for Car 2 R equals = nothing mi​ / 10 gal ​(Type an integer or decimal rounded to one decimal place as​ needed.) Sample standard deviation for Car 1 s equals = nothing mi​ / 10 gal ​(Type an integer or decimal rounded to one decimal place as​ needed.) Sample standard deviation for Car 2

sequals=nothing

mi​ / 10 gal

​(Type an integer or decimal rounded to one decimal place as​ needed.)

Which car would the customer buy and​ why?

A.

Car

22​,

because it has a lower mean gas mileage.

B.

Car

11​,

because it has a larger range of gas mileage.

C.

Car

22​,

because it has a lower sample standard​ deviation, hence more predictable gas mileage.

D.

There is very little difference between the two cars.
Answer #1

(a)

Sample mean of Car 1 (1) is got as follows:

mi​ / 10 gal

(b)

Sample mean of Car 2 (2) is got as follows:

mi​ / 10 gal

(c)

Median of Car 1 (1) is got as follows:

Arranging numbers in ascending order, we get:

159,164,208,208,216,235,239,251,263,283,285,297,298,305,307

n = 15

Median = (15+1)/2th item = 8th item = 251

So,

Median of Car 1 (1) = 251 mi​ / 10 gal

(d)

Median of Car 2 (2) is got as follows:

Arranging numbers in ascending order, we get:

180,205,215,221,232,241,243,246,246,248,259,261,272,294,298

n = 15

Median = (15+1)/2th item = 8th item = 246

So,

Median of Car 2 (2) = 246 mi​ / 10 gal

(e)

Range for Car 1 = Maximum - Minimum = 307 - 159 = 148 mi​ / 10 gal

(f)

Range for Car 2 = Maximum - Minimum = 298 - 180 = 118 mi​ / 10 gal

(g)

x (x - ) (x - )2
235 -12.8667 165.5511
208 -39.8667 1589.3511
239 -8.8667 78.6178
216 -31.8667 1015.4844
208 -39.8667 1589.3511
297 49.1333 2414.0844
285 37.1333 1378.8844
159 88.8667 7897.2844
298 50.1333 2513.3511
251 3.1333 9.8178
164 - 83.8667 7033.6178
307 59.1333 3496.7511
283 35.1333 1234.3511
305 57.1333 3264.2178
263 15.1333 229.0178
Total = 33909.7333

Sample standard deviation for Car 1 =s1 is given by:

s1 =    mi​ / 10 gal

(h)

x (x - ) (x - )2
215 - 29.0667 844.8711
205 -39.0667 1526.2044
232 -12.0667 145.6044
221 -23.0667 532.0711
246 1.9333 3.7378
248 3.9333 15.4711
241 -3.0667 9.4044
272 27.9333 780.2711
243 -1.0667 1.1378
298 53.9333 2908.8044
180 -64.0667 4104.5378
246 1.9333 3.7378
259 14.9333 223.0044
294 49.9333 2493.3378
261 16.9333 286.7378
Total = 13878.9333

Sample standard deviation for Car 2 =s2 is given by:

s2 = mi​ / 10 gal

(i)

Correct option:

C.

Car 2 because it has a lower sample standard​ deviation, hence more predictable gas mileage.